# CliCARE: Grounding Large Language Models in Clinical Guidelines for Decision Support over Longitudinal Cancer Electronic Health Records

### v1

@@ 3.2 Trajectory-Guideline Alignment
Our guideline knowledge graph, $G_{g}$, is a normative, static graph formalized as $G_{g}=(E_{g},R_{g})$. It is constructed based on authoritative CPGs (e.g., NCCN), where $E_{g}$ represents abstract medical concepts such as Cancer, ClinicalSituation, and Treatment. The edges, $R_{g}$, represent logical and recommendation relationships, collectively forming a graph that represents an idealized clinical workflow. We first perform a basic entity-level alignment between the patient’s temporal graph, $G_{t}$, and $G_{g}$ using medical ontologies. Subsequently, the clinical history for each patient $p$ is organized from $G_{t}$ into a time-ordered sequence of Clinical Events, resulting in a Temporal Trajectory $Tr_{p}=\langle e_{1},e_{2},\dots,e_{m}\rangle$. The set of all patient trajectories is denoted as $\{Tr_{p}\}_{p=1}^{P}$. Concurrently, we systematically enumerate all possible normative treatment workflows from $G_{g}$ to form a set of paths $\{Pa_{k}\}_{k=1}^{K}$, where each path $Pa_{k}=\langle s_{1},s_{2},\dots,s_{l}\rangle$ represents a recommended sequence of guideline steps.
@@
We developed a global matching strategy based on deep semantic representations to directly assess the similarity between the entire patient trajectory and each candidate guideline path. Specifically, the overall matching score for a candidate path $Pa_{k}=\langle s_{1},s_{2},\dots,s_{l}\rangle$ with the patient trajectory $Tr_{p}$ is computed as follows. First, for each step $s_{j}$ in the guideline path, we identify the most semantically similar event in the patient’s complete event trajectory $Tr_{p}$. This similarity is calculated using a BERT model, $f_{\text{BERT}}$, pre-trained on biomedical text. Finally, we aggregate the best-match similarities for all steps to derive the total score for the path. Formally, the matching score is defined as:
@@
[…]
@@
The aforementioned method generates a ranked list of candidate alignment paths for each patient trajectory. However, purely algorithmic matching may fail to capture the complexities of clinical logic. Therefore, we introduce an LLM as a Clinical Reasoner, $f_{LLM}$, to perform reranking. We provide the LLM with a rich-context prompt in a zero-shot manner, which includes the patient’s trajectory $Tr_{p}$, the top-N candidate normative paths $\{Pa_{1},\dots,Pa_{N}\}$, and their corresponding matching scores $\{\text{Score}_{1},\dots,\text{Score}_{N}\}$. The LLM’s task is to evaluate which candidate alignment is the most clinically plausible and to output a reranked list: $\langle Pa^{\prime}_{1},\dots,Pa^{\prime}_{N}\rangle=f_{LLM}(Tr_{p},\{\langle Pa_{k},\text{Score}_{k}\rangle\}_{k=1}^{N})$.
@@ 4.1 Evaluation Method
To assess the quality of generated text, we developed a Human-validated LLM-as-a-Judge component. This component assesses two primary tasks: retrospective Clinical Summary, referred to as $T_{\textbf{CS}}$, and prospective Clinical Recommendation, referred to as $T_{\textbf{CR}}$. Our methodology employs a concise, four-dimensional rubric, co-designed with senior oncologists (see Appendix A for details), to assess Factual Accuracy, Completeness & Thoroughness, Clinical Soundness, and Actionability & Relevance. A robust LLM judge is prompted to assign a score ranging from 1 (poor) to 5 (excellent) for each dimension.
To address the known systematic biases of LLM judges, including positional bias, verbosity, and self-enhancement (Zheng et al. 2023; Wang et al. 2024), and after verifying their presence in our specific context (Appendix E), we implement a robust two-part mitigation protocol. First, to ensure rating stability and reduce model bias, we create a judging ensemble composed of three powerful LLMs—GPT-4.1, Claude 4.0 Sonnet, and Gemini 2.5 Pro—using their averaged score. Second, to eliminate ordering effects, all items are presented in a randomly shuffled sequence during the evaluation.
@@
where $n$ is the number of samples and $d_{i}$ is the difference in ranks for each sample. A high correlation provides strong evidence that our LLM judge functions as a reliable and scalable proxy for human expert assessment. This justifies its application for large-scale evaluations throughout our experiments.
@@ 4.2 Datasets
We evaluated our framework using two large-scale clinical datasets, with detailed analyses provided in Appendices B and C. The first dataset is a private Chinese collection, referred to as CancerEHR. It contains longitudinal records for 2,000 patients from Liaoning Cancer Hospital. These records span extensive periods—some exceeding two decades—resulting in inputs of up to 20,000 tokens. The dataset includes a variety of data types, such as physicians’ orders, laboratory results, and surgical notes. The second dataset is derived from the publicly available MIMIC-IV dataset (Johnson et al. 2023), filtered to include only patients with cancer-related diagnoses, which we refer to as MIMIC-Cancer. This dataset provides a focus on disease progression similar to CancerEHR; however, the language and data structure differ, offering a robust test of our method’s generalizability. For brevity, in the subsequent implementation and results sections, we will refer to the two datasets as $D_{\textbf{CEHR}}$ for CancerEHR and $D_{\textbf{MC}}$ for MIMIC-Cancer. Similarly, the two primary tasks will be abbreviated as $T_{\textbf{CS}}$ for Clinical Summary and $T_{\textbf{CR}}$ for Clinical Recommendation.
@@ 4.3 Baselines
We compare our proposed CliCARE framework against a variety of robust baseline methods. These include standard retrieval-augmented generation (RAG) pipelines implemented with powerful open-source models such as Mistral-7B and its instruction-tuned variant (Jiang et al. 2023), Qwen3-8B (Yang et al. 2025), and the domain-specific BioMistral-7B (Labrak et al. 2024). Additionally, we evaluate more advanced KG-enhanced RAG techniques designed for long-context or knowledge-intensive tasks. The selected methods include BriefContext (Zhang et al. 2024), which employs a Map-Reduce strategy, as well as several Graph-Aware RAG frameworks, such as GNN-RAG (Feng et al. 2024), KG2RAG (e Shi et al. 2024), and the healthcare-focused MedRAG (Zhao et al. 2025).
@@ 4.4 Implementation Details
In the knowledge graph alignment stage, the threshold is set to 0.7 when using BERT to calculate semantic (cosine) similarity in the initial step. During the fine-tuning stage, we divided the 2,000-sample dataset into a training set of 1,800 samples and a test set of 200 samples, with 10% of the training data reserved for validation. The key hyperparameters for training include a batch size of 1, a maximum context length of 20,000 tokens, and an initial learning rate of $5\times 10^{-5}$ with a cosine scheduler. We utilized BF16 for mixed-precision training, set the maximum output length to 4,096 tokens, and trained for 3 epochs. All experiments were conducted using a configuration of four NVIDIA A800 GPUs.
@@
Table 1: CliCARE Outperforms RAG Baselines on Clinical Generation Tasks. Scores are assigned by our Human-validated LLM-as-a-Judge. The asterisk (*) denotes KG-enhanced RAG variants.
@@ 4.5 Experimental Results
Acknowledging the limitations of traditional metrics for clinical tasks, we validated our LLM-as-a-Judge protocol against three experienced oncologists. To ensure a feasible yet representative assessment, we created a validation subset by randomly sampling generated outputs from eight different models. Our protocol minimized bias by evaluating these outputs column-wise and presenting the Clinical Summary and Recommendation tasks together for a comprehensive review (details in Appendix D). The results demonstrate a strong positive correlation between the automated ratings and those of the experts. Specifically, the Spearman’s rank correlation ($\rho$) between our LLM judge’s scores and the physicians’ mean scores was approximately 0.7, confirming that our metric serves as a reliable proxy for human expert judgment.
@@
As demonstrated in Table 2, our framework’s knowledge structuring offers a significant advantage over standard RAG. The performance uplift is most pronounced on the complex CancerEHR dataset, where nearly all models exhibit substantial gains. Notably, the improvements for Qwen-3-8B and Deepseek-R1 are the largest in their respective groups, with their Clinical Summary scores increasing by a remarkable +1.688 and +2.279, respectively. This underscores that even advanced models require a coherent structure for effective reasoning on complex records. On the simpler MIMIC-Cancer dataset, while the absolute gains are smaller, CliCARE still delivers a distinct and consistent advantage. For instance, it elevates the score of a strong baseline like GPT-4.1 from 4.419 to 4.737, a gain of +0.318. While the uplift is nearly universal, we do note a single case of minor performance degradation, confirming the intricate nature of these tasks.
@@
Further analysis of record length reveals distinct performance patterns, as detailed in Table 4. The smaller model, Qwen-3-8B, performs optimally on Short length records but exhibits a significant decline in quality when processing the longest records, particularly with the complex CancerEHR data. In contrast, the more powerful Gemini-2.5-Pro model demonstrates strong and consistent performance across all record lengths. Notably, when guided by the CliCARE framework, it achieves its highest scores on the longest record segments for both datasets. This finding suggests that CliCARE effectively organizes extensive clinical histories enables advanced models to leverage richer context for enhanced reasoning.
@@ 5 Conclusion
We introduced CliCARE, a framework for reliable clinical decision support that transforms cancer EHRs into Temporal Knowledge Graphs and aligns them with clinical guidelines. This approach addresses key challenges in long-context reasoning and hallucination, enabling both small specialist and large generalist models to significantly outperform strong baselines. We also validated a robust LLM-as-a-Judge protocol that correlates highly with expert oncologist assessments, representing a significant advancement toward deploying trustworthy AI in clinical practice.
@@ Appendix A Human Evaluation Questionnaire and Protocol
  • Anonymized outputs from a random sample of eight models—including both locally deployed (CliPAGE_8B,etc.) and API-based (CliPAGE,etc.) systems—were compared against each other and the gold-standard labels.
@@ Appendix G Implementation Details of the Alignment Method
1:Clinical Guidelines $D_{g}$; Patient Clinical Data $D_{p}$; Pre-trained BERT model $f_{\text{BERT}}$; Large Language Model $f_{\text{LLM}}$; Bootstrapping iterations $I$; Confidence threshold $\theta$; Weighting factor $\alpha$ for scoring.
@@
5:Utilize $f_{\text{LLM}}$ to extract a static $KG$ from $D_{g}$ and a Temporal $TKG$ from $D_{p}$.
@@ Appendix H PROMPT OF Answer Label GENERATION
As shown in Figure 8, this is a sample prompt for the CliPAGE method. The prompt is structured into five primary components: Longitudinal Cancer EHRs, Current record, Retrieval-Augmented Generation (RAG) content, a section for the Clinical Summary, and a section for the Clinical Recommendation. These elements combine to form a single, comprehensive prompt that guides the model’s response generation process.
Figure 8: A Sample Prompt for the CliPAGE Method Highlighting Its Key Components.

### v2

@@ 3.2 Trajectory-Guideline Alignment
Our guideline knowledge graph, $G_{g}$, is a normative, static graph formalized as $G_{g}=(E_{g},R_{g})$. It is constructed based on authoritative CPGs, where $E_{g}$ represents abstract medical concepts such as Cancer, ClinicalSituation, and Treatment. The edges, $R_{g}$, represent logical and recommendation relationships, collectively forming a graph that represents an idealized clinical workflow. We first perform a basic entity-level alignment between the patient’s temporal graph, $G_{t}$, and $G_{g}$ using medical ontologies. Subsequently, the clinical history for each patient $p$ is organized from $G_{t}$ into a time-ordered sequence of Clinical Events, resulting in a Temporal Trajectory $Tr_{p}=\langle e_{1},e_{2},\dots,e_{m}\rangle$. The set of all patient trajectories is denoted as $\{Tr_{p}\}_{p=1}^{P}$. Concurrently, we systematically enumerate all possible normative treatment workflows from $G_{g}$ to form a set of paths $\{Pa_{k}\}_{k=1}^{K}$, where each path $Pa_{k}=\langle s_{1},s_{2},\dots,s_{l}\rangle$ represents a recommended sequence of guideline steps.
@@
We developed a global matching strategy based on deep semantic representations to directly assess the similarity between the entire patient trajectory and each candidate guideline path. Specifically, the matching score for a candidate path $Pa_{k}=\langle s_{1},s_{2},\dots,s_{l}\rangle$ with the patient trajectory $Tr_{p}$ is computed as follows. First, for each step $s_{j}$ in the guideline path, we identify the most semantically similar event in the patient’s event trajectory $Tr_{p}$. This similarity is calculated using a BERT model, $f_{\text{BERT}}$, pre-trained on biomedical text. Finally, we aggregate the best-match similarities for all steps to derive the total score for the path. Formally, the matching score is defined as:
@@
where the $\text{desc}(\cdot)$ function retrieves the text description of a node, and cos_sim computes the cosine similarity between two vectors. A higher score indicates a better alignment between the normative path $Pa_{k}$ and the patient’s experience. The optimal matching path $Pa^{*}$ is then determined by selecting the candidate with the highest score:
@@
The method described above generates a ranked list of candidate alignment paths for each patient trajectory. However, purely algorithmic matching may fail to capture the complexities of clinical logic. Therefore, we introduce an LLM as a Clinical Reasoner, $f_{LLM}$, to perform reranking. We provide the LLM with a rich-context prompt in a zero-shot manner, which includes the patient’s trajectory $Tr_{p}$, the top-N candidate normative paths $\{Pa_{1},\dots,Pa_{N}\}$, and their corresponding matching scores $\{\text{Score}_{1},\dots,\text{Score}_{N}\}$. The LLM’s task is to evaluate which candidate alignment is the most clinically plausible and to output a reranked list: $\langle Pa^{\prime}_{1},\dots,Pa^{\prime}_{N}\rangle=f_{LLM}(Tr_{p},\{\langle Pa_{k},\text{Score}_{k}\rangle\}_{k=1}^{N})$.
@@ 4.1 Evaluation Method
To assess the quality of generated text, we developed an Expert-Validated LLM-as-a-Judge component. This component assesses two primary tasks: retrospective Clinical Summary, referred to as $T_{\textbf{CS}}$, and prospective Clinical Recommendation, referred to as $T_{\textbf{CR}}$. Our methodology employs a concise, four-dimensional rubric, co-designed with senior oncologists, to assess Factual Accuracy, Completeness & Thoroughness, Clinical Soundness, and Actionability & Relevance. The LLM judge is prompted to assign a score ranging from 1 (poor) to 5 (excellent) for each dimension.
To address the systematic biases of LLM judges, including positional bias, verbosity, and self-enhancement (Zheng et al. 2023; Wang et al. 2024), and after verifying their presence in our specific context, we implement a robust two-part mitigation protocol. First, to ensure rating stability and reduce model bias, we create a judging ensemble composed of three powerful LLMs, such as GPT-4.1 (Achiam et al. 2023), Claude 4.0 Sonnet (Anthropic 2025), and Gemini 2.5 Pro (Comanici et al. 2025), using their averaged score. Second, to eliminate ordering effects, all items are presented in a randomly shuffled sequence during the evaluation.
@@
where $n$ is the number of samples and $d_{i}$ is the difference in ranks for each sample. A high correlation provides evidence that our LLM judge functions as a reliable and scalable proxy for expert assessment. This justifies its application for large-scale evaluations throughout our experiments.
@@ 4.2 Datasets
We evaluated our framework using two large-scale clinical datasets. The first dataset is a private Chinese collection, referred to as CancerEHR. It contains longitudinal records for 2,000 patients from Liaoning Cancer Hospital. These records span extensive periods—some exceeding two decades—resulting in inputs of up to 20,000 tokens. The dataset includes a variety of data types, such as physicians’ orders, laboratory results, and surgical notes. The second dataset is derived from the publicly available MIMIC-IV dataset (Johnson et al. 2023), filtered to include only patients with cancer-related diagnoses, which we refer to as MIMIC-Cancer. This dataset provides a focus on disease progression similar to CancerEHR; however, the language and data structure differ, offering a robust test of our method’s generalizability. For brevity, in the subsequent implementation and results sections, we will refer to the two datasets as $D_{\textbf{CEHR}}$ for CancerEHR and $D_{\textbf{MC}}$ for MIMIC-Cancer. Similarly, the two primary tasks will be abbreviated as $T_{\textbf{CS}}$ for Clinical Summary and $T_{\textbf{CR}}$ for Clinical Recommendation.
@@ 4.3 Baselines
We compare our proposed CliCARE framework against a variety of robust baseline methods. These include standard RAG pipelines implemented with powerful open-source models such as Mistral-7B and its instruction-tuned variant (Jiang et al. 2023), Qwen3-8B (Yang et al. 2025a), and the domain-specific BioMistral-7B (Labrak et al. 2024). Additionally, we evaluate more advanced KG-enhanced RAG techniques designed for long-context or knowledge-intensive tasks. The selected methods include BriefContext (Zhang et al. 2025), which employs a Map-Reduce strategy, as well as several Graph-Aware RAG frameworks, such as GNN-RAG (Mavromatis and Karypis 2025), KG2RAG (Zhu et al. 2025), and the healthcare-focused MedRAG (Zhao et al. 2025). We evaluate these RAG frameworks using both the open-source models and leading generalist models, including Deepseek-R1 (Guo et al. 2025), Gemini 2.5 Pro, GPT-4.1, and Claude-4.0-Sonnet.
@@ 4.4 Implementation Details
In the knowledge graph alignment stage, the threshold is set to 0.7 when using BERT to calculate semantic cosine similarity in the initial step. During the fine-tuning stage, we divided the 2,000-sample dataset into a training set of 1,800 samples and a test set of 200 samples, with 10% of the training data reserved for validation. The key hyperparameters for training include a batch size of 1, a maximum context length of 20,000 tokens, and an initial learning rate of $5\times 10^{-5}$ with a cosine scheduler. We utilized BF16 for mixed-precision training, set the maximum output length to 4,096 tokens, and trained for 3 epochs. All experiments were conducted using a configuration of 4 NVIDIA A800 GPUs.
@@
Table 1: CliCARE Outperforms RAG Baselines on Clinical Generation Tasks. Scores are assigned by our Expert-Validated LLM-as-a-Judge. The asterisk (*) denotes KG-enhanced RAG variants.
@@ 4.5 Experimental Results
Acknowledging the limitations of traditional metrics for clinical tasks, we validated our LLM-as-a-Judge protocol against three experienced oncologists. To ensure a feasible yet representative assessment, we created a validation subset by randomly sampling generated outputs from eight different models. Our protocol minimized bias by evaluating these outputs column-wise and presenting the Clinical Summary and Recommendation tasks together for a comprehensive review. The results demonstrate a strong positive correlation between the automated ratings and those of the experts. Specifically, the Spearman’s rank correlation $\rho$ between our LLM judge’s scores and the physicians’ mean scores was approximately 0.7, confirming that our metric serves as a reliable proxy for human expert judgment.
@@
As demonstrated in Table 2, our framework’s knowledge structuring offers a significant advantage over standard RAG. The performance uplift is most pronounced on the complex CancerEHR dataset, where nearly all models exhibit substantial gains. Notably, the improvements for Qwen-3-8B and Deepseek-R1 are the largest in their respective groups, with their Clinical Summary scores increasing by +1.688 and +2.279, respectively. This underscores that even advanced models require a coherent structure for effective reasoning on complex records. On the simpler MIMIC-Cancer dataset, while the absolute gains are smaller, CliCARE still delivers a distinct and consistent advantage. For instance, it elevates the score of a strong baseline like GPT-4.1 from 4.419 to 4.737, a gain of +0.318. While the uplift is nearly universal, we do note a single case of minor performance degradation, confirming the intricate nature of these tasks.
@@
Further analysis of record length reveals distinct performance patterns, as detailed in Table 4. The smaller model, Qwen-3-8B, performs optimally on Short length records but exhibits a decline in quality when processing the longest records, particularly with the complex CancerEHR data. In contrast, the more powerful Gemini-2.5-Pro model demonstrates strong and consistent performance across all record lengths. Notably, when guided by the CliCARE framework, it achieves its highest scores on the longest record segments for both datasets. This suggests that CliCARE effectively organizes extensive clinical histories and enables advanced models to leverage richer context for enhanced reasoning.
@@ 5 Conclusion
We introduced CliCARE, a framework for reliable clinical decision support that transforms cancer EHRs into Temporal Knowledge Graphs and aligns them with clinical guidelines. This approach addresses key challenges in long-context reasoning and hallucination, enabling both small specialist and large generalist models to significantly outperform baselines. We also validated a robust LLM-as-a-Judge protocol that correlates highly with expert oncologist assessments, representing a significant advancement toward deploying trustworthy AI in clinical practice. Future work should focus on generalizing this framework to a wider range of clinical domains.
Acknowledgments
This work was supported by the Natural Science Foundation of Liaoning Province, China (No. 2025-MS-036) and the Doctoral Student Scientific Research Innovation Project of Northeastern University (No. 02190022125036).
@@ Appendix A Human Evaluation Questionnaire and Protocol
  • Anonymized outputs from a random sample of eight models—including both locally deployed (CliCARE_8B,etc.) and API-based (CliCARE,etc.) systems—were compared against each other and the gold-standard labels.
@@ Appendix G Implementation Details of the Alignment Method
1:Clinical Guidelines $D_{g}$; Patient Clinical Data $D_{p}$; Pre-trained BERT model $f_{\text{BERT}}$; Large Language Model $f_{\text{LLM}}$; Bootstrapping iterations II; Confidence threshold $\theta$; Weighting factor $\alpha$ for scoring.
@@
5:Utilize $f_{\text{LLM}}$ to extract a static KGKG from $D_{g}$ and a Temporal TKGTKG from $D_{p}$.
@@ Appendix H PROMPT OF Answer Label GENERATION
As shown in Figure 8, this is a sample prompt for the CliCARE method. The prompt is structured into five primary components: Longitudinal Cancer EHRs, Current record, Retrieval-Augmented Generation (RAG) content, a section for the Clinical Summary, and a section for the Clinical Recommendation. These elements combine to form a single, comprehensive prompt that guides the model’s response generation process.
Figure 8: A Sample Prompt for the CliCARE Method Highlighting Its Key Components.